# SCORE: Story Coherence and Retrieval Enhancement for AI Narratives

### v2

@@ Abstract
Large Language Models (LLMs) can generate creative and engaging narratives from user-specified input, but maintaining coherence and emotional depth throughout these AI-generated stories remains a challenge. In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies. By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF Ramos et al. (2003) and cosine similarity Rahutomo et al. (2012) to identify related episodes and enhance the overall story structure. Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
1 Introduction
@@
Deep learning has transformed multiple domains including NLP, time series analysis and computer vision Qiu et al. (2025a, b, 2024); Li et al. (2024b); Li and Deng (2023); Li et al. (2024a). Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  Tao et al. (2024). However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge McAdams (2006). For instance, Khatun and Brown (2024) pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
@@
Similarly, Liu et al. (2024) highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement. These challenges indicate a need for more structured approaches in narrative generation that can better manage character arcs, plot developments, and emotional progression throughout the story.
@@
In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agents. Zhang et al. (2024) conducted a comprehensive survey on these mechanisms, identifying effective memory designs that help mitigate inconsistencies in narrative development—a challenge common to both interactive agents and narrative generation tasks. Additionally, Park et al. (2023) introduced generative agents that simulate human-like behavior using memory modules. These agents track the state of a wide array of interactable objects in a sandbox environment, ensuring consistent reasoning and enabling the smooth functioning of a simulated society. These researches inspired the design of our new framework.
@@
In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) Lewis et al. (2020), which dynamically incorporates relevant context to enhance narrative coherence. Expanding on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements. Our key contributions are:
@@
  • We outperform baseline GPT model Radford et al. (2019) in detecting continuity errors and maintaining overall narrative coherence.
2 Method
@@
2.1 Continuity Analysis and Key Item Status Correction
@@
2.2 Key Item Interaction Analysis
@@
2.3 Similarity-Based Episode Evaluation and Sentiment Analysis
We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries. It begins by loading summaries, full episode content, and key item states from structured JSON files. The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS Douze et al. (2024) and OpenAI embeddings. This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
@@ 2.3 Similarity-Based Episode Evaluation and Sentiment Analysis
3 Experiments
3.1 Dataset Preparation
3.2 Baselines
Table 1: Performance comparison of models with and without SCORE
@@ 3.2 Baselines
3.3 Evaluation Process
@@ Appendix B Theoretical Analysis
B.1 Proof of the Continuity Analysis of Critical Item State Correction
@@
We define the narrative entropy function of a state sequence by McAdams (2006). Let
@@
B.2 Theory for complex question performance gain
B.2.1 Preliminaries
@@
The details and underlying proof are provided in the section B.2.
@@
Table 2: Evaluation Metrics Results
@@
Table 3: Evaluation of Episode 26 by GPT4o and use our prompt
@@
Table 4: Evaluation of Episode 26 in GPT4o playground
@@
Table 5: Evaluation of Episode 26 by GPT4o-mini and use our prompt
@@
Table 6: Evaluation of Episode 26 in GPT4o-mini playground

### v3

@@ Abstract
Large Language Models (LLMs) can generate creative and engaging narratives from user-specified input, but maintaining coherence and emotional depth throughout these AI-generated stories remains a challenge. In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies. By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF [1] and cosine similarity [2] to identify related episodes and enhance the overall story structure. Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
Index Terms:
Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human Feedback
I Introduction
@@
Deep learning has transformed multiple domains including NLP, time series analysis and computer vision [3, 4, 5, 6, 7, 8]. Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  [9]. However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge [10]. For instance, [11] pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
@@
Similarly, [12] highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement. These challenges indicate a need for more structured approaches in narrative generation that can better manage character arcs, plot developments, and emotional progression throughout the story.
@@
In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agents. [13] conducted a comprehensive survey on these mechanisms, identifying effective memory designs that help mitigate inconsistencies in narrative development—a challenge common to both interactive agents and narrative generation tasks. Additionally, [14] introduced generative agents that simulate human-like behavior using memory modules. These agents track the state of a wide array of interactable objects in a sandbox environment, ensuring consistent reasoning and enabling the smooth functioning of a simulated society. These researches inspired the design of our new framework.
@@
In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) [15], which dynamically incorporates relevant context to enhance narrative coherence. Expanding on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements. Our key contributions are:
@@
  • We outperform baseline GPT model [16] in detecting continuity errors and maintaining overall narrative coherence.
II Method
@@
II-A Continuity Analysis and Key Item Status Correction
@@
II-B Key Item Interaction Analysis
@@
II-C Similarity-Based Episode Evaluation and Sentiment Analysis
We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries. It begins by loading summaries, full episode content, and key item states from structured JSON files. The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS [17] and OpenAI embeddings. This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
@@ II-C Similarity-Based Episode Evaluation and Sentiment Analysis
TABLE I: Performance comparison of models with and without SCORE
@@
III Experiments
III-A Dataset Preparation
III-B Baselines
III-C Evaluation Process
@@ Appendix B Theoretical Analysis
B-A Proof of the Continuity Analysis of Critical Item State Correction
@@
We define the narrative entropy function of a state sequence by [10]. Let
@@
B-B Theory for complex question performance gain
B-B1 Preliminaries
@@
The details and underlying proof are provided in the section B-B.
@@
TABLE II: Evaluation Metrics Results
@@
TABLE III: Evaluation of Episode 26 by GPT4o and use our prompt
@@
TABLE IV: Evaluation of Episode 26 in GPT4o playground
@@
TABLE V: Evaluation of Episode 26 by GPT4o-mini and use our prompt
@@
TABLE VI: Evaluation of Episode 26 in GPT4o-mini playground